A) Each month, a manufacturer stocks one warehouse with good x and two warehouses with good y. Monthly sales for x and y (as fractions of a full warehouse) are then random, described by the following joint density:

T1,9) 1-k(x+y), if x € [0, 1], y € (0.2) 10. else

Find k to make f(x, y) a legitimate density function.

B) For the joint density given in part B, find Cov(X, Y ) and Corr(X, Y ).

IMPORTANT: If you're going to write please make sure your writing is neat/easy to read
Answer #1

2533333333335.23333 3 33333333.5.2 Ans Given fix, y) = $ 1-K(x+y) 2€ [1], y €[0,1] else This is a joint density function in x299 - *[ <1+ Beste) du] =1 2- Klit 2017] = 1 - 3K = -1 Hence the value of not so pff becomes f(x,y) = 1 - ļlity) 2€6,1) YE [0for this we have to find for & freys (marginal densities we have 00 fxcol= | fax,y) dy 4 - Fyry) = 1 fingsde Here ferrys = diSsanlines Simila - (4) Jar) IyM [13+ [ ] - La - ) - 9 - Llo- ) + ) +1 : (- 2 부 - 름 글 : ( So fly = E(X)= (x forces de - (-) dxMy = ey)- , yfylgidy 4=6 = 547-5)do = 1 19) - 4 (8) Bo Cov(x,y) = cov (X,Y Try ferry) de dy - Mx My 53 +3/-1-3 109) desdy – (19) - 9146 - 2 - [19)+ (9)] – 25 31- $ [ +4 ] - 28 F1-8-20 Covex,y) = { - -0.01234 cov wkt cor (x,y)=9= cover, y) 6x Gy for wЕс”) - в w 5м . се ( 3) - **(-3)250 - eets/-5), «да 9 Яхь — = 5 9хь L- hz = (2) - (*), з | 1 - 1 1 1 •r (4-x).) сан . В19) - 9146 - 2 - [19)+ (9)] – 25 31- $ [ +4 ] - 28 F1-8-20 Covex,y) = { - -0.01234 cov wkt cor (x,y)=9= cover, y) 6x Gy for w